a particle starts from t=0 with a velocity of 8j and moves inthe xy plane with a constant acceleraton of 4i+2j......at the moment when its x coordinate is 29m,what is its speed?y coordinate?

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shivanshi

 ux = 0

uy= 8 m/sec

ax = 4 m/sec2

ay = 2 m/sec2


 use formula  Sx= ux t + 1/2 ax t2

                      29 = 0 + 1/2 *4 *t2

                     t= √(29/2)  sec

   Vx = Ux  + ax t

        = 4√(29/2)

 Vy =Uv + ay t

      =8 + 2√(29/2)

speed  =[Vx2 +Vy2]1/2

y co ordinate Sy=Uy t + 1/2 ayt2

                            =8√(29/2)  +29/2

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin